OpenStudy (anonymous):

Use logarithmic differentiation to calculate y prime....Please show work y= ((x-1)^5) * (x+2)^(1/4))/((x^2+7)^(1/2))

OpenStudy (anonymous):

is it like this? \[y=\frac{(x-1)^5(x+2)^\frac{1}{4}}{(x^2+7)^\frac{1}{2}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first you want to take ln of both sides: \[\ln y = \ln \left(\frac{(x−1)^5(x+2)^\frac{1}{4}}{(x^2+7)^\frac{1}{2}}\right)\] then using properties of ln, such as:\[\ln (a\cdot b)=\ln a + \ln b\]and\[\ln \left(\frac{a}{b}\right)=\ln a - \ln b\]and\[\ln x^a = a \ln x\]re write the right side as sums/differences/products to make it easier to take the derivative of.